This problem is really difficult, I just can't seem to figure it out. Will someone PLEASE HELP ME!
If:
z = 6(cos 205 degrees + i sin 205 degrees) and w = 3(cos 85 degrees + i sin 85 degrees), which of the following is (z/w) in polar form?
1. 2(cos 290 degrees + i sin 290 degrees)
2. 0.5(cos(-120 degrees) + i sin (-120 degrees))
3. 18(cos 290 degrees + i sin 290 degrees)
4. 2(cos 120 degrees + i sin 120 degrees)
Thanks again for all of your help, I really appreciate it!
Natalie Answer by Edwin McCravy(20081) 
 This problem is really difficult, I just can't seem to figure it out. Will someone PLEASE HELP ME!
If:
z = 6(cos 205 degrees + i sin 205 degrees) and w = 3(cos 85 degrees + i sin 85 degrees), which of the following is (z/w) in polar form?
1. 2(cos 290 degrees + i sin 290 degrees)
2. 0.5(cos(-120 degrees) + i sin (-120 degrees))
3. 18(cos 290 degrees + i sin 290 degrees)
4. 2(cos 120 degrees + i sin 120 degrees)
Thanks again for all of your help, I really appreciate it!
Natalie
You just have to learn the rules for multiplying and dividing complex
numbers in trigonometric (or polar) form:
The rule for multiplying complex numbers in trig (polar) form is:
"multiply the r's and add the angles":
You don't need that one now but you will on other problems.
--------------------------------------
The rule for dividing complex numbers in polar form is similar:
"Divide the r's and subtract the angles":
z = 6(cos 205 degrees + i sin 205 degrees)
w = 3(cos 85 degrees + i sin 85 degrees)
Choice 4.
Edwin